# Inferring CSM Properties of Type II SNe Using a Magnitude-Limited ZTF Sample

### v2

@@ 6.3 CSM Mass and Radial Extent
For events with lower inferred $M_{CSM}$ ($<10^{-2.5}$ $\textup{M}_{\odot}$), the impact of CSM on the early light curve is too weak to constrain $R_{CSM}$. While many of these SNe are likely to possess non-negligible CSM, material at larger radii would likely not influence the early evolution, and would be too diffuse to influence the later light curve evolution in a clearly discernible way unless the mass is extremely high (e.g. Irani et al., 2024).
A limitation of the approach from this work is that for the fastest rising events ($\leq$ 1 d), which likely require dense, compact CSM to achieve such rapid evolution, our measured rise times would represent upper limits, unresolved fast rises might require more substantial $M_{CSM}$ or closer and more compact CSM. While individual CSM parameters may have uncertainties due to model assumptions, observational constraints and scatter present in relationships, this frequency of substantial CSM in our volume corrected sample represents a robust statistical result, independent of the precise CSM parameterisation.
6.4 Implications on Mass-Loss Mechanisms
With a minimum $M_{CSM}$, $10^{-2.5}$ $\textup{M}_{\odot}$, and observationally supported fiducial values for $v_{wind}$ and $R_{CSM}$ of 10 km s-1 and $5\times 10^{14}$ cm, respectively, we find a characteristic $\dot{M}$ of 2×$\times 10^{-4}$ $\textup{M}_{\odot}$ yr-1. This characteristic value is higher, by $\sim$ two orders of magnitude, than values inferred from observations of local group RSGs e.g. $\dot{M}\sim 10^{-6}$ $\textup{M}_{\odot}$ yr-1 (e.g. Vink et al., 2001; Smith, 2014; Beasor & Davies, 2018; Beasor et al., 2020).
@@ 6.4 Implications on Mass-Loss Mechanisms
Given a $v_{wind}$ of 10 km s-1 and the minimum and maximum values of $R_{CSM}$, $10^{14}$ – $10^{15}$ cm, we calculate the range of timescales in which the material is removed to be 3 – 32 yrs. The distinctly short timescales of mass-loss, compared to the lifespan of this evolutionary stage, further supports the need for a period of ‘enhanced’ mass-loss. We argue the higher rates of mass-loss estimated from early photometry found here and by other studies (e.g. Irani et al., 2024; Jacobson-Galán et al., 2024b; Silva-Farfán et al., 2024) are probing the end-of-life mass-loss rather than the typical $\dot{M}$  of RSGs.
@@ 7 Conclusions
This supports findings in recent literature that CSM interactions contribute significantly to the early light curve and are prevalent in a large set Type II SNe. While not ubiquitous across Type II SNe, we show possessing large amounts of CSM in common amongst Type II SNe progenitors. For the $\sim$ 36% of Type II SNe (excluding Types IIn and IIb SNe) where CSM interactions dominate, we find that dense CSM both shortens the rise time to peak luminosity and enhances the early-time brightness. We have further highlighted the need to reconcile and address the disparity between light curve derived $\dot{M}$ values and $\dot{M}$ from RSG observations – which are typically larger by $\approx 2$ orders of magnitudes than the rates inferred from local group RSG observations (e.g. van Loon et al., 2005; Mauron & Josselin, 2011; Smith, 2014; Beasor et al., 2020; Stroh et al., 2021; Strotjohann et al., 2023).
@@ 9 Data Availability
The light curves and empirical properties measured for the SNe, as they relate to this work, can be found in the supplementary material.
@@ A.5 M23 Luminosity-Rise
Fig. 18 shows the distribution of $t_{25\_75}$ vs. $M_{g,peak}$ for the theoretical light curve grid from M23. We present 10,000 model points using the same $V_{max}\times M_{ZAMS}^{-2.35}$ weighting scheme applied in Fig. 8, which accounts for both observational selection effects and the IMF (e.g. Salpeter, 1955). The top panel colour-codes data by $\dot{M}$, while the bottom panel uses $R_{CSM}$.
This visualisation reveals how CSM properties strongly influence the distribution of SNe in the $t_{25\_75}$-$M_{g,peak}$ plane. Fast risers ($t_{25\_75}$leq 5 d) typically have confined, dense CSM characterised by higher $\dot{M}$ and smaller $R_{CSM}$, producing moderately more luminous peaks. In contrast, slower risers ($t_{25\_75}$ $>$ 5 d) typically exhibit less confined and less dense CSM with lower overall $M_{CSM}$ values and larger $R_{CSM}$. Notably, even within the slower-rising population, the most luminous events still require substantial CSM masses, confirming that CSM mass remains a fundamental driver of peak luminosity across the distribution. The clear separation between these populations emerges naturally from the underlying physics rather than from arbitrary parameter choices, suggesting fundamental differences in mass-loss mechanisms or progenitor structures. This bi-modality provides valuable context for interpreting the observed distribution of Type II SNe in our sample.
@@ A.6 $M_{Fe,Core}$ Measurements
We extend the predictive capabilities of multi-output GPR extend to estimating the iron core mass, $M_{Fe,Core}$, of the progenitor through Eq. 3, which exploits a tight correlation between $M_{Fe,Core}$ and the plateau luminosity at 50 d in simulated Type IIP light curves (e.g. Barker et al., 2022; Barker et al., 2023). The theoretical correlation indicates that more massive stellar cores lead to more energetic and luminous SNe, notably enhancing the bolometric luminosity during the plateau phase at approximately 50 d post-explosion (Barker et al., 2022). To quantify this relationship, we utilise the bolometric plateau luminosity at 50 d, $L_{bol,50d}$. The plateau length is measured by analysing the gradient along the light curve and identifying significant changes in the slope. The ZTF $g$ and $r$ band magnitudes are measured at 50 d after the plateau onset. A bolometric correction is then applied to convert these magnitudes into bolometric luminosity – we adopt the methodology described by Lyman et al. (2014).
@@
While this correlation provides a useful estimate of the core mass, it assumes a direct relationship between core mass and explosion energy that, in reality, may be complicated by ejecta properties (mass and H-richness). Higher ejecta masses or more H-richness can extend and diminish the plateau luminosity independent of core mass (e.g. Goldberg, 2022). The use of luminosity at 50 d may be particularly sensitive to hydrogen envelope mass variations, as it assumes complete H retention (e.g. Goldberg, 2022; Fang et al., 2025). A more robust approach might utilise the luminosity at half the plateau duration, which better accounts for diversity in envelope masses and better isolates the core mass contribution to the light curve evolution (Fang et al., 2025).
Figure 19: $M_{Fe,Core}$ KDE distribution.
Figure 20: Type II KDE for $M_{Fe,Core}$ along with the associated 80% CI. The weighted distribution is in dark green (dashed line) and the unweighted normalised histogram is in black. .
@@ A.9 M23 $M_{CSM}$ Lower Limit
Figure 22: Dependence of $M_{g,peak}$ on $M_{CSM}$ under different parameter variations, with fixed Ni mass and explosion energy. Top left: Varying $\dot{M}$ with fixed $\beta$ and $R_{CSM}$. Top right: Varying $\beta$ with fixed $\dot{M}$ and $R_{CSM}$. Bottom left: Varying $R_{CSM}$ with fixed $\dot{M}$ and $\beta$. Bottom right: Combined variation of all CSM parameters ($\dot{M}$, $\beta$, $R_{CSM}$). Each panel shows results for different progenitor masses (10 – 18 $\textup{M}_{\odot}$). Below $M_{CSM}$ $\approx 10^{-2.5}$ $\textup{M}_{\odot}$, CSM properties do not significantly influence the peak magnitude, indicating a transition to CSM-negligible evolution.

### v3

@@ 6.3 CSM Mass and Radial Extent
For events with lower inferred $M_{CSM}$ ($<10^{-2.5}$ $\textup{M}_{\odot}$), the impact of CSM on the early light curve is too weak to constrain $R_{CSM}$. While many of these SNe are likely to possess non-negligible CSM, material at larger radii would likely not influence the early evolution, and would be too diffuse to influence the later light curve evolution in a clearly discernible way unless the mass is extremely high (e.g., Irani et al., 2024).
A limitation of the approach from this work is that for the fastest rising events ($\leq$ 1 d), which likely require dense, compact CSM to achieve such rapid evolution, our measured rise times would represent upper limits, unresolved fast rises might require more substantial $M_{CSM}$ or closer and more compact CSM. While individual CSM parameters may have uncertainties due to model assumptions, observational constraints and scatter present in relationships, this frequency of substantial CSM in our volume corrected sample represents a robust statistical result, independent of the precise CSM parameterisation. Our analysis is further constrained by confidence in classifications, an issue we explore in greater detail in section A.10.
6.4 Implications for Mass-Loss Mechanisms
With a minimum $M_{CSM}$, $10^{-2.5}$ $\textup{M}_{\odot}$, and observationally supported fiducial values for $v_{wind}$ and $R_{CSM}$ of 10 km s-1 and $5\times 10^{14}$ cm, respectively, we find a characteristic $\dot{M}$ of 2×$\times 10^{-4}$ $\textup{M}_{\odot}$ yr-1. This characteristic value is higher, by $\sim$ two orders of magnitude, than values inferred from observations of local group RSGs e.g., $\dot{M}\sim 10^{-6}$ $\textup{M}_{\odot}$ yr-1 (e.g., Vink et al., 2001; Smith, 2014; Beasor & Davies, 2018; Beasor et al., 2020).
@@ 6.4 Implications for Mass-Loss Mechanisms
Given a $v_{wind}$ of 10 km s-1 and the minimum and maximum values of $R_{CSM}$, $10^{14}$ – $10^{15}$ cm, we calculate the range of timescales in which the material is removed to be 3 – 32 yrs. The distinctly short timescales of mass-loss, compared to the lifespan of this evolutionary stage, further supports the need for a period of ‘enhanced’ mass-loss. We argue the higher rates of mass-loss estimated from early photometry found here and by other studies (e.g., Irani et al., 2024; Jacobson-Galán et al., 2024b; Silva-Farfán et al., 2024) are probing the end-of-life mass-loss rather than the typical $\dot{M}$  of RSGs.
@@
As an alternative to ‘enhanced’ $\dot{M}$ in the centuries before core-collapse, the dense chromosphere model of Fuller & Tsuna (2024) offers a compelling explanation for the apparent disparity between observed RSG mass-loss rates and those inferred from early SN evolution (see Fuller & Tsuna, 2024). A chromosphere exists in approximate hydrostatic equilibrium, with significantly higher densities above the stellar surface than predicted by constant or $\beta$-law wind velocity models, despite maintaining $\dot{M}$ more consistent with local group RSG measurements (see Fig. 4 in Fuller & Tsuna, 2024). This model is able to naturally produce the rapid photometric and spectroscopic evolution observed in fast-rising SNe while preserving realistic progenitor properties. While our $M_{CSM}$ estimates would remain largely unaffected, the inferred $\dot{M}$ would decrease substantially due to the significantly different velocity structure in Fuller & Tsuna (2024) compared to those used in M23. Thus, the chromosphere model represents a promising alternative to ‘enhanced’ $\dot{M}$, potentially resolving a tension between $\dot{M}$ measured from local RSGs and early SN evolution, though additional development is required to fully understand the impact of chromospheres on the photometric and spectroscopic evolution across diverse SN populations.
@@ 7 Conclusions
This supports findings in recent literature that CSM interactions contribute significantly to the early light curve and are prevalent in a large set Type II SNe. While not ubiquitous across Type II SNe, we show possessing large amounts of CSM in common amongst Type II SNe progenitors. For the $\sim$ 36% of Type II SNe (excluding Types IIn and IIb SNe) where CSM interactions dominate, we find that dense CSM both shortens the rise time to peak luminosity and enhances the early-time brightness. We have further highlighted the need to reconcile and address the disparity between light curve derived $\dot{M}$ values and $\dot{M}$ from RSG observations – which are typically larger by $\approx 2$ orders of magnitudes than the rates inferred from local group RSG observations (e.g., van Loon et al., 2005; Mauron & Josselin, 2011; Smith, 2014; Beasor et al., 2020; Stroh et al., 2021; Strotjohann et al., 2024).
@@ 9 Data Availability
The light curves and empirical properties measured for the SNe, as they relate to this work, can be found here: [https://doi.org/10.5281/zenodo.15229515].
@@ A.1 Heavily Host-Extinguished
Table 8 contains significantly dust-extinguished Type II SNe we identify in our sample. These events are characterised by distinctly red colours, $(g-r)_{g,max}$$\geq$ 0.25 mag, at peak and moderate rise times ,$t_{25\_75}$leq20 d, placing them in a unique region of parameter space as illustrated in Fig. 2. We correct for host extinction using $(g-r)_{g,max}$ and apply this correction to these SNe only, as described in Section section 2.4.
@@ A.2 Rise Time Recovery
To assess the impact of the sampling function on our GPR measurements, we conducted a systematic resampling experiment using well-sampled light curves with well-constrained $t_{25\_75}$ values. This involved taking thoroughly observed events (e.g., ZTF18aacnlxz/SN 2020aavr) and resampling their light curves (simulating alternative sampling functions) to match the observation cadence of more sparsely observed light curves in our sample. Figure 13 illustrates the resulting distribution of measured rise times across different intrinsic $t_{25\_75}$ values.
Our analysis demonstrates that the GPR process reliably distinguishes between fast-rising ($t_{25\_75}$leq 5 d) and slower-rising ($t_{25\_75}$ $>$ 5 d) events. For the fastest risers ($t_{25\_75}$ between 1–2 d), we observe substantial uncertainty with a spread of $\approx$ 0.7 dex. This improves to $\approx$ 0.4 dex for moderate risers ($t_{25\_75}$ between 3–5 d) and further to $\approx$ 0.2 dex for slower-rising events ($t_{25\_75}$ $>$ 5 d).
@@ A.3 Sample Redshift Distribution
In Fig. reffig:RedshfitDist, we show the distribution of redshift, $z$, across our SN sample. The upper panel shows the $z$ distribution for our full dataset, and the lower panel displays the highly complete ($\sim$95%) sample limited to events with peak apparent magnitudes $m_{peak}$leq 18.5 mag.
@@ A.4 Peak Colours
In Figs. 15(a)17(b), we present the KDE distributions (left) and ECDFs (right) of peak $g-r$ colours ($(g-r)_{g,max}$) for Type II, Type IIn and Type IIb SNe. The panels display distributions for standard Type II (top), Type IIn (middle), and Type IIb (bottom) SNe. We show various statistical quantities for each distribution in Table 4.
@@ A.5 M23 Luminosity-Rise
Fig. 18 shows the distribution of $t_{25\_75}$ vs. $M_{g,peak}$ for the theoretical light curve grid from M23. We present 10,000 model points using the same $V_{max}\times M_{ZAMS}^{-2.35}$ weighting scheme applied in Fig. 8, which accounts for both observational selection effects and the IMF (e.g., Salpeter, 1955). The top panel colour-codes data by $\dot{M}$, while the bottom panel uses $R_{CSM}$.
This visualisation reveals how CSM properties strongly influence the distribution of SNe in the $t_{25\_75}$-$M_{g,peak}$ plane. Fast risers ($t_{25\_75}$leq 5 d) typically have confined, dense CSM characterised by higher $\dot{M}$ and smaller $R_{CSM}$, producing moderately more luminous peaks. In contrast, slower risers ($t_{25\_75}$ $>$ 5 d) typically exhibit less confined and less dense CSM with lower overall $M_{CSM}$ values and larger $R_{CSM}$. Notably, even within the slower-rising population, the most luminous events still require substantial CSM masses, confirming that CSM mass remains a fundamental driver of peak luminosity across the distribution.
The clear separation between these populations emerges naturally from the underlying physics rather than from arbitrary parameter choices, suggesting fundamental differences in mass-loss mechanisms or progenitor structures. This bi-modality provides valuable context for interpreting the observed distribution of Type II SNe in our sample.
@@ A.6 $M_{Fe,Core}$ Measurements
We extend the predictive capabilities of multi-output GPR extend to estimating the iron core mass, $M_{Fe,Core}$, of the progenitor through Eq. 3, which exploits a tight correlation between $M_{Fe,Core}$ and the plateau luminosity at 50 d in simulated Type IIP light curves (e.g., Barker et al., 2022; Barker et al., 2023). The theoretical correlation indicates that more massive stellar cores lead to more energetic and luminous SNe, notably enhancing the bolometric luminosity during the plateau phase at approximately 50 d post-explosion (Barker et al., 2022). To quantify this relationship, we utilise the bolometric plateau luminosity at 50 d, $L_{bol,50d}$. The plateau length is measured by analysing the gradient along the light curve and identifying significant changes in the slope. The ZTF $g$ and $r$ band magnitudes are measured at 50 d after the plateau onset. A bolometric correction is then applied to convert these magnitudes into bolometric luminosity – we adopt the methodology described by Lyman et al. (2014).
@@
While this correlation provides a useful estimate of the core mass, it assumes a direct relationship between core mass and explosion energy that, in reality, may be complicated by ejecta properties (mass and H-richness). Higher ejecta masses or more H-richness can extend and diminish the plateau luminosity independent of core mass (e.g., Goldberg, 2022). The use of luminosity at 50 d may be particularly sensitive to hydrogen envelope mass variations, as it assumes complete H retention (e.g., Goldberg, 2022; Fang et al., 2025). A more robust approach might utilise the luminosity at half the plateau duration, which better accounts for diversity in envelope masses and better isolates the core mass contribution to the light curve evolution (Fang et al., 2025).
Figure 19: Type II KDE for $M_{Fe,Core}$ along with the associated 80% CI. The weighted distribution is in dark green (dashed line) and the unweighted normalised histogram is in black.
@@ A.8 M23 Radial Extent Predictions
Fig. 20 shows the multivariate analysis comparing our polynomial regression-predicted CSM radial extent values (y-axis) against the corresponding M23 model values (x-axis), following an approach similar to Figure 10. This systematic evaluation examines the performance of polynomial fits across different degrees (1st, 2nd, and 3rd order) and specific parameter regimes to determine the optimal method for characterising this relationship. Unlike our $M_{CSM}$ analysis, this investigation of $R_{CSM}$ is restricted to M23 models with $M_{CSM}$geq 1×$\times 10^{-2.5}$ $\textup{M}_{\odot}$, which Section section 5.1 identifies as the threshold above which CSM significantly influences both $M_{g,peak}$ and $t_{25\_75}$. The results demonstrate that $R_{CSM}$ can only be reliably constrained for events with substantial $M_{CSM}$.
@@ A.9 M23 $M_{CSM}$ Lower Limit
Figure 21: Dependence of $M_{g,peak}$ on $M_{CSM}$ under different parameter variations, with fixed Ni mass and explosion energy. Top left – Varying $\dot{M}$ with fixed $\beta$ and $R_{CSM}$. Top right – Varying $\beta$ with fixed $\dot{M}$ and $R_{CSM}$. Bottom left – Varying $R_{CSM}$ with fixed $\dot{M}$ and $\beta$. Bottom right – Combined variation of all CSM parameters ($\dot{M}$, $\beta$, $R_{CSM}$). Each panel shows results for different progenitor masses (10 – 18 $\textup{M}_{\odot}$). Below $M_{CSM}$ $\approx 10^{-2.5}$ $\textup{M}_{\odot}$, CSM properties do not significantly influence the peak magnitude, indicating a transition to CSM-negligible evolution.
@@
A.10 Impact of Systematic Misclassifications
Most SN classifications from the BTS rely on the low-resolution SEDM spectrograph (R $\sim$100). The limited spectral resolution and typically single-epoch observations near maximum light can make distinguishing certain SN subclasses challenging, particularly Type IIb from Type II and, to a lesser extent, Type IIn from Type II or host emission. Consequently, our Type II sample might contain some level of contamination from misclassified events, a consideration we quantitatively address here.
To quantify potential classification biases, we conducted Kolmogorov-Smirnov (KS) tests comparing Type II and Type IIb populations. KS tests of the unweighted $M_{g,peak}$ and $t_{25\_75}$ distributions yielded $p$-values of 0.038 and 0.0030, respectively, indicating statistically significant differences between these populations. We identified an approximately 5% shortfall of Type IIb SNe in our sample ($7.22^{+2.40}_{-1.84}$versus the expected $\sim$12.5from Shivvers et al. 2017’s volume-complete sample). To assess the impact of possible misclassifications, we applied a conservative approach by removing the fastest-rising 5% of Type II SNe – those most likely to be misclassified Type IIbs and have the largest impact on our results – and recalculated the $M_{CSM}$ KDE distribution. The fraction of Type II SNe with $M_{CSM}$$\geq 10^{-2.5}$ $\textup{M}_{\odot}$ remained consistent (38 – 41%) with our original finding ($\sim$36%). This represents the most extreme scenario, confirming that potential misclassifications affect our results by less than 1 $\sigma$.
Similarly, our analysis yields a Type IIn rate of $4.34^{+1.49}_{-1.09}\%$ relative to Type II SNe, consistent with Shivvers et al. (2017). We consider the possibility of misclassification between regular Type II SNe and Type IIn events to be minimal for several reasons: (1) Type IIn SNe typically exhibit higher luminosities and represent a small fraction of the overall population, resulting in negligible statistical impact after $V_{max}$ weighting; (2) BTS routinely conducts follow-up observations using higher-resolution spectrographs for suspected Type IIn events to refine classification; and (3) wee see that $<$2% of our Type IIn sample exhibits photometric characteristics resembling typical Type II events (e.g., $t_{25\_75}$$\leq$ 3 d and $M_{g,peak}$$>-18$ mag, which represent the median values for our unweighted Type II sample; Table 4). Our robust classification methodology ensures complete Type IIn identification, and our focus on CSM around fast-rising SNe means the longer evolution timescales of Type IIn events minimally impact our conclusions.
We have also considered the potential impact of peculiar events resembling SN 1987A on our results. Such objects, characterised by moderate peak luminosities combined with unusually slow rise times (e.g., ZTF18acbwaxk), represent rare occurrences in the local universe. If several such events were misclassified or included within our sample, their statistical contribution would remain minimal given our $V_{max}$ weighting and large sample size.